Information on Bowgie Chemicals direct materials costs follows.

 





Actual quantities of direct materials used
30,800
Actual costs of direct materials used$137,800
Standard price per unit of direct materials$4.28
Flexible budget for direct materials$121,300

 

Bowgie Chemicals has no materials inventories.

 

Required:
a. 
Prepare a short report for management showing Bowgie Chemicals’s direct materials price and efficiency variances.
b. (Appendix) Prepare the journal entries to record the purchase and use of the direct materials using standard costing.
Answer #1

   Actual quantities of direct materials used                              30,800

 Actual costs of direct materials used                                    $137,800 

Standard price per unit of direct materials                               $4.28 

Flexible budget for direct materials                                        $121,300 


A) 137,800/ 30,800= 4.474025974

( 4.474025974- 4.28)* 30,800= 5,975.9999992 or 5,976 U DM  price variance


137,800- 121,300= 16,500- 5,976= 10, 524 U DM efficiency variance


B)

WIP inventories                             121,300 debit

materials efficiency variance            5,976 debit

materials price variance                  10,524 debit

  accounts payable                                                137,800 credit